OpenStudy (anonymous):

A mole of carbon dioxide (CO2) has 12.0 grams of carbon (C) and 32.0 grams of oxygen (O). What is the percent composition of carbon in carbon dioxide?

OpenStudy (hlilly2413):

Step 1: Find the molar mass of CO2. To do that: Molar Mass \[CO _{2}\] = 1(mass of 1 carbon atom found on the periodic table) + 2(mass of one oxygen atom from the periodic table). Step 2: to find the percentage of carbon in CO2 you would do: \[\frac{ 1 \times mass of 1 carbon atom from periodic table }{ Molar Mass of CO2 } \times 100\] Step 3: Do the same thing with oxygen, but instead of multiplying the top mass by 1, you multiply it by 2 because there are two oxygen atoms in the molecule. *Honestly, you can skip step 2 and just do step 3 since that is all your question is asking. However, your percentages from step 2 and step 3 should equal 100. Hope it helps!

OpenStudy (hlilly2413):

Oops I lied, you can skip step 3 since you are looking for carbon, but if you calculate it- both percentages should equal 100.

OpenStudy (hlilly2413):

Also, your units are g/mol except for the final result which is of course a percentage.